OpenStudy (anonymous):

What is the volume of a sphere that has a diameter of 36 yd? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (jdoe0001):

\(\Large \textit{volume of a sphere}=\cfrac{4}{3}\pi r^3\qquad r=radius=\frac{diameter}{2}\)